1. Two charges are on the x-axis. Charge q1 = +35 nC is at x = 3.00 cm while charge q2 = -65nC is at x = 6.00 cm. Find the



magnitude and direction of



a) the electric force that q1 exerts on q2; and



b) the electric force the q2 exerts on q1.



2. A positive charge of 10-8 coulombs (C) experiences a force of 0.95 N when located at a certain point in an electric field.



Find the intensity (magnitude) of the electric field at that point.

1
Expert's answer
2022-03-06T18:06:27-0500

1.

"F_1=F_2=\frac{kq_1q_2}{(x_1-x_2)^2}=22.8~mN,"

2.

"E=\frac Fq=95~\frac{MN}C."